OpenStudy (anonymous):

I need help, no answers just help! Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. How would i start this problem?

jimthompson5910 (jim_thompson5910):

what is the purchase price

OpenStudy (anonymous):

$0.50 per flower plus 40%

jimthompson5910 (jim_thompson5910):

that's the sale price

jimthompson5910 (jim_thompson5910):

it costs them $0.50 to buy each flower they then turn around and sell each flower at $0.50 + 40% of $0.50 = 0.50 + 0.4*0.5 = 0.70 dollars

jimthompson5910 (jim_thompson5910):

so if they buy x flowers it will cost them a total of 0.50x dollars let's make C the cost function, so C = 0.50x

jimthompson5910 (jim_thompson5910):

and if they sell all x flowers bought at $0.70, then the revenue function is R = 0.70x

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the profit, call it P, is equal to the difference of the revenue and cost Profit = Revenue - Cost P = R - C P = 0.70x - 0.50x P = (0.70 - 0.50)x P = (0.20)x P = 0.20x

jimthompson5910 (jim_thompson5910):

so if you buy and sell x flowers, then your profit is 0.20x dollars

jimthompson5910 (jim_thompson5910):

you want the profit to be $500, which means P = 0.20x 500 = 0.20x ... plug in P = 500 then solve for x

jimthompson5910 (jim_thompson5910):

I'll let you do this

OpenStudy (anonymous):

x=2,500 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you must buy and sell 2500 flowers to get a profit of $500

OpenStudy (anonymous):

Thank you so much !

jimthompson5910 (jim_thompson5910):

you're welcome